# Characterization of Multiple Channels Room Temperature Readout Electronics for Large Transition-Edge Sensor Array

### v1

@@ Abstract
Transition-edge sensor (TES) is a highly sensitive device that is capable of detecting extremely low levels of energy. It is characterised by low noise performance and high energy resolution. A mature method for reading out TES signal is through time-division multiplexing (TDM) direct current superconducting quantum interference device (SQUID). In a TDM system, the signal readout chain represents a significant source of noise other than the TES intrinsic noise. The noise generated by TES is in the range of several tens to several hundreds of $pA/\sqrt{Hz}$. In order to ensure the high energy resolution of TES, it is necessary to ensure that the noise contribution from the room temperature readout electronics is less than $10$ $pA/\sqrt{Hz}$ above 100 $Hz$. In this work, we have designed a low-noise, high-resolution room temperature readout circuit for TDM. The equivalent current noise contribution of ADC is about $0.05$ $pA/\sqrt{Hz}$ above 100 $Hz$ and $0.46$ $pA\sqrt{Hz}$ under 30:1 multiplexing. The resolution of the analog to digital converter (ADC) is larger than 11.5 bits, which can reconstruct the TES siganl without distortion. The readout board, which has eight channels, has JESD204B serial ports, which has greatly simplified the space of room temperature electronics. The readout chain is based on multi-threaded CPU processing and can transfer data at 2 $Gbps$ for each channel in real time. This readout board can be used in a TDM system with smaller size for large TES arrays.
@@ 1 Introduction
TES micro-calorimeters and bolometers have ultra low thermal noise and ultra high energy resolution. Therefore TESs are widely used in particle physics experiments, such as X-ray [1], $\gamma$-ray [2], dark matter [3], neutrinos [4], cosmic microwave background  [5] detection and so on. For example, in X-ray detection experiments, the energy resolution is on the order of a few electronvolts ($eV$)  [6]. In axion or dark photon detection, the energy resolution is on the order of sub-$eV$ [3, 7]. One way to readout the TES signal is based on a direct current SQUID [8]. The signal is fully amplified by SQUID at cryogenic refrigerator, which provides enhanced immunity to interference and thus allows for matching with the room temperature readout circuit.
@@ 2.2 Testing system
Although there are two-stage SQUIDs in TDM system, only one SQUID array is enough for verifying the functionality of the readout circuit and analyze the noise contribution of ADC. The SQUID array was tested in a pulse turbo refrigerator at 4 $K$ temperature, as illustrated in Figure 3. The SQUID array comprises 100 SQUID units in series, which can provide further signal amplification. To prevent ground loops, the SQUID ground was connected to the room temperature electronics ground at a single point.
@@ 3.1 SQUID flux response curve and noise analysis
The system bandwidth becomes smaller after TDM multiplexing, which in turn gives rise to the aliasing of Nyquist noise [12]. In order to avoid the crosstalk caused by TDM, the row selection time $t_{row}$ for each line of TES array to be read out can be set as $t_{row}=2{\pi}{\tau}_{OL}$ [18]. ${\tau}_{OL}$ is the time constant relative to the open-loop bandwidth ${f}_{OL}=1/(2{\pi}{\tau}_{OL})$ when we model the TDM as a single-pole, low-pass filter. Thus the integrated effective noise bandwidth of TDM is ${f}_{eff}=1/(4{\tau}_{OL})$. The Nyquist bandwidth of TDM is ${f}_{Nyq}=1/(2N_{rows}t_{row})$. After considering ${f}_{eff}$ and ${f}_{Nyq}$, we can get the relationship between the number of multiplexed rows $N_{rows}$ and the system noise should be satisfied:
$S_{I,MUX}\begin{cases}=S_{I,ADC}\sqrt{\frac{f_{eff}}{f_{Nyq}}}=S_{I,ADC}\sqrt{{\pi}N_{rows}}&,N_{rows}>1\\ =S_{I,ADC}&,N_{rows}=1\end{cases}$ (3.2)
$S_{I,MUX}$ denotes the total current noise contribution of the ADC after multiplexing. $S_{I,ADC}$ denotes the single-pixel ADC current noise contribution. The system noise is illustrated in Figure 5.
Figure 5: ADC noise contribution analysis. The red line represents the 10 $pA/\sqrt{Hz}$ noise density refer to input coil of input SQUID. The blue line depicts the ADC noise contribution of the 8th channel, which was measured by inserting a $50$ $\Omega$ terminal at the ADC input. The green line shows the 30:1 multiplexed noise contribution of ADC.
@@
30:1 TDM current noise contribution ($pA\sqrt{Hz}$) $0.87@100Hz$ $0.39@100kHz$ $0.68@100Hz$ $0.29@100kHz$ $0.49@100Hz$ $0.19@100kHz$ $0.46@100Hz$ $0.19@100kHz$

### v2

@@ Abstract
Transition-edge sensor (TES) is a highly sensitive device that is capable of detecting extremely low levels of energy. It is characterised by low noise performance and high energy resolution. A mature method for reading out TES signal is through time-division multiplexing (TDM) direct current superconducting quantum interference device (SQUID). In a TDM system, the signal readout chain represents a significant source of noise other than the TES intrinsic noise. The noise generated by TES is in the range of several tens to several hundreds of $pA/\sqrt{Hz}$. In order to ensure the high energy resolution of TES, it is necessary to ensure that the noise contribution from the room temperature readout electronics is less than $10$ $pA/\sqrt{Hz}$ above 100 $Hz$. In this work, we have designed a low-noise, high-resolution room temperature readout circuit for TDM. The equivalent current noise contribution of ADC is about $0.05$ $pA/\sqrt{Hz}$ above 100 $Hz$ and $0.46$ $pA\sqrt{Hz}$ under 30:1 multiplexing. The resolution of the analog to digital converter (ADC) is larger than 11.5 bits, which can reconstruct the TES signal without distortion. The readout board, which has eight channels, has JESD204B serial ports, which has greatly simplified the space of room temperature electronics. The readout chain is based on multi-threaded CPU processing and can transfer data at 2 $Gbps$ for each channel in real time. This readout board can be used in a TDM system with smaller size for large TES arrays.
@@ 1 Introduction
TES micro-calorimeters and bolometers have ultra low thermal noise and ultra high energy resolution. Therefore TESs are widely used in particle physics experiments, such as X-ray [1], $\gamma$-ray [2], dark matter [3], neutrinos [4], cosmic microwave background  [5] detection and so on. For example, in X-ray detection experiments, the energy resolution is on the order of a few electronvolts ($eV$)  [1, 6]. In axion or dark photon detection, the energy resolution is on the order of sub-$eV$ [3, 7]. One way to readout the TES signal is based on a direct current SQUID [8]. The signal is fully amplified by SQUID at cryogenic refrigerator, which provides enhanced immunity to interference and thus allows for matching with the room temperature readout circuit.
@@ 2.2 Testing system
Although there are two-stage SQUIDs in TDM system, only one SQUID array is enough for verifying the functionality of the readout circuit and analyze the noise contribution of ADC. The SQUID array was tested in a pulse tubo refrigerator at 4 $K$ temperature, as illustrated in Figure 3. The SQUID array comprises 100 SQUID units in series, which can provide further signal amplification. To prevent ground loops, the SQUID ground was connected to the room temperature electronics ground at a single point.
@@ 3.1 SQUID flux response curve and noise analysis
The system bandwidth becomes smaller after TDM multiplexing, which in turn gives rise to the aliasing of Nyquist white noise [12]. In order to avoid the crosstalk caused by TDM, the row selection time $t_{row}$ for each line of TES array to be read out can be set as $t_{row}=2{\pi}{\tau}_{OL}$ [18]. ${\tau}_{OL}$ is the time constant relative to the open-loop bandwidth ${f}_{OL}=1/(2{\pi}{\tau}_{OL})$ when we model the TDM as a single-pole, low-pass filter. Thus the integrated effective noise bandwidth of TDM is ${f}_{eff}=1/(4{\tau}_{OL})$. The Nyquist bandwidth of TDM is ${f}_{Nyq}=1/(2N_{rows}t_{row})$. After considering ${f}_{eff}$ and ${f}_{Nyq}$, we can get the relationship between the number of multiplexed rows $N_{rows}$ and the aliased white noise should be satisfied:
$S_{I,MUX,W}\begin{cases}=S_{I,ADC,W}\sqrt{\frac{f_{eff}}{f_{Nyq}}}=S_{I,ADC,W}\sqrt{{\pi}N_{rows}}&,N_{rows}>1\\ =S_{I,ADC,W}&,N_{rows}=1\end{cases}$ (3.2)
$S_{I,MUX,W}$ denotes the total aliased current white noise contribution of the ADC after multiplexing. $S_{I,ADC,W}$ denotes the single-pixel ADC current white noise contribution. The total system noise $S_{I,MUX}=S_{I,ADC}+S_{I,MUX,W}-S_{I,ADC,W}$ when $N_{rows}>1$ is illustrated in Figure 5. $S_{I,ADC}$ denotes the single-pixel ADC total current noise contribution, which inclues the non-linear $1/f$ noise and white noise $S_{I,ADC,W}$.
Figure 5: ADC noise contribution analysis. The red line represents the 10 $pA/\sqrt{Hz}$ noise density refer to input coil of input SQUID. The blue line depicts the ADC current noise contribution $S_{I,ADC}$ of the 8th channel, which was measured by inserting a $50$ $\Omega$ terminal at the ADC input. The green line shows the 30:1 multiplexed noise contribution $S_{I,MUX}$ of ADC.
@@
30:1 TDM current noise contribution ($pA\sqrt{Hz}$) $0.28@100Hz$ $0.23@100kHz$ $0.26@100Hz$ $0.22@100kHz$ $0.24@100Hz$ $0.21@100kHz$ $0.24@100Hz$ $0.21@100kHz$